OpenStudy (anonymous):

What is the value of m in the following equation: 3(2m - 9 + 6m) = 3m +29 -7m? -3 2 5 120

OpenStudy (anonymous):

6m - 27 + 18m = 3m + 29 - 7m 24m - 27 = -4m + 29 20m - 27 = 29 20m = 56 m = 2 4/5

OpenStudy (haseeb96):

3(8m-9)=-4m+29 24m-27=-4m+29 24m+4m=29+27 28m=56 m=56\28 m=2

OpenStudy (anonymous):

Sorry, Idk how I got that, it's probably wrong

OpenStudy (haseeb96):

hence B is the Correct answer

OpenStudy (anonymous):

Thanks @Haseeb96 and @Baseballa101